OpenStudy (anonymous):

find x and y intercepts of... y-x^2=7x ??

OpenStudy (anonymous):

Same, put 0 for y to get x intercept, 0 for x to get y intercept

OpenStudy (anonymous):

i did... 0-x^2=7x i just dont know whee to go from there!

OpenStudy (anonymous):

Move -x^2 to the other side 0=x^2+7x now factor

OpenStudy (anonymous):

0-x^2=7x +x^2 +x^2 0=x^2+7x x(x+7) x=0 x=7 ????

OpenStudy (anonymous):

Itd be x=0 and x+7=0 so x=0 and x=-7

OpenStudy (anonymous):

oh ok i got it:)

OpenStudy (anonymous):

what about y=x^3-7x^2+6x as gonna use the quadic formula?

OpenStudy (anonymous):

Are you still trying to find x and y intercepts?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

y=x^3-7x^2+6x = x(x^2 -7x +6) = x(x-6)(x-1) the rest is same as the above process.

OpenStudy (anonymous):

so x= 0, 6 ,1 ???

OpenStudy (anonymous):

Yep

OpenStudy (anonymous):

for y-intercept, yes.

OpenStudy (anonymous):

Actually that'd be x intercepts.

OpenStudy (anonymous):

o ya, lol

OpenStudy (anonymous):

yea x intercepts :)

OpenStudy (anonymous):

im imbarrased to ask again lol what about one like y= x^3 / 3x+5 ??

OpenStudy (anonymous):

Same idea, x intercept would be x^3/3x+5=0

OpenStudy (anonymous):

from there?

OpenStudy (anonymous):

Multiply both sides by 3x+5 x^3=0 x=0

OpenStudy (anonymous):

3x+5=x^3 ?

OpenStudy (anonymous):

Nope, 3x+5*0=?

OpenStudy (anonymous):

i wanna know the steps this is just homework but i got a test in a week i want to know howto do it

OpenStudy (anonymous):

Alright, you're dividing by 3x+5 on the left side right? so to undo that you do the opposite operation to both sides...multiplication by 3x+5...it eliminates it from the denominator and 3x+5*0=0 So x^3=0

OpenStudy (anonymous):

o ok 0 but you still got the x^3 dont you>

OpenStudy (anonymous):

Yep, take the cube root of both sides.... x=0

OpenStudy (anonymous):

o......ok thanks again for your help im gona go try and do some of my own ill probably be back in a while if i get stuck on another... thanls :)

OpenStudy (anonymous):

you're welcome